OpenStudy (savannah_noelle):

The length of the longer leg of a right triangle is 6inches more than twice the length of the shorter leg. The length of the hypotenuse is 9inches more than twice the length of the shorter leg. Find the side lengths of the triangle.

satellite73 (satellite73):

we need a variable, what do you pick?

OpenStudy (savannah_noelle):

x

satellite73 (satellite73):

ok then if the shorter leg is x, what is the longer leg? i.e how would you write "6 more than twice x"?

OpenStudy (savannah_noelle):

6+2x?

satellite73 (satellite73):

yes

satellite73 (satellite73):

and how would you write 9 more twice x?

OpenStudy (savannah_noelle):

9+2x

satellite73 (satellite73):

yes now pythgoras

satellite73 (satellite73):

\[a^2+b^2=c^2\]so in your case \[x^2+(2x+6)^2=(2x+9)^2\]

satellite73 (satellite73):

multiply that mess out, combine like terms, and solve the resulting quadratic equation you will get two solutions, ignore the negative one

OpenStudy (savannah_noelle):

Okay so 15?

satellite73 (satellite73):

yes

satellite73 (satellite73):

well the short one is \(x=15\) the other two you find by replacing \(x\) by \(15\)

OpenStudy (savannah_noelle):

Thank you so much!

satellite73 (satellite73):

yw solved that quadratic quick huh?

OpenStudy (savannah_noelle):

Longer is 36 right?

satellite73 (satellite73):

yes

OpenStudy (savannah_noelle):

and the hypotenuse is 39?

satellite73 (satellite73):

yes

satellite73 (satellite73):

you have \(15,36,39\) if you divide all by 3 you get the famous \(5,12,13\) right triangle

OpenStudy (savannah_noelle):

Thanks again! :)